OpenStudy (anonymous):

I need helpwith these 3 questions please! 1. radical 10 (11+ radical 5) 2. 7 over 10i 3. radical 49 over (3+4i)-(2-5i) Please help!

OpenStudy (anonymous):

What are the questions here? I'm not sure I understand. Do you need to simplify? Rationalize?

OpenStudy (anonymous):

yes simplify. im sorry

OpenStudy (anonymous):

1. 10√(1/5) - √45 - 12√(5/9) 10√(1/5) - √9√5 -(12/3)√5 10√(1/5) - 3√5 - 4√5 10√(1/5) - 7√5 2. You have to know that each numbers has factors 10 has the factors 10 and 1 and 5 and 2 12 has factors 12 and 1, 6 and 2 and 4 and 3 You simplify by doing the same to numerator (top number) and denomiator (bottom number). The only common factors are 1 and 2. Since 1 leaves you with the same expression you take the higher number 2 to get 5/6. If you knew multiplication tables you would see the answer inherently. I am not sure those tables are taught in schools anymore. They are here: http://www.mathsisfun.com/timestable.htm … 3.i dont know

OpenStudy (anonymous):

\[\frac{7}{10i}=\frac{7}{10i}\times \frac{i}{i}=\frac{7i}{-10}=-\frac{7i}{10}\]

OpenStudy (anonymous):

#3, first of \[\sqrt{49}=7\] and \[(3+4i)-(2-5i)=1-i\]so start with \[\frac{7}{1-i}\]

OpenStudy (anonymous):

then \[\frac{7}{1-i}=\frac{7}{1-i}\times \frac{1+i}{1+i}=\frac{7(1+i)}{2}\]

OpenStudy (anonymous):

last line because \[(a+bi)(a-bi)=a^2+b^2\]

OpenStudy (anonymous):

then standard form looks like \[\frac{7}{2}+\frac{7}{2}i\]

OpenStudy (anonymous):

the answers to the first question is A.-11 radical 10 + 5 radical 2 B. -5 radical 2 + 11i radical 10 C. 5 radical 2 - 11i radical 10 D. 11 radical 10 + 5 radical 2

OpenStudy (anonymous):

yeah, they were all wrong..

OpenStudy (anonymous):

thanks for your effort though